OpenStudy (anonymous):

In parallelogram DEFG, DH = x + 5, HF = 2y, GH = 3x – 1, and HE = 5y + 4. Find the values of x and y.

OpenStudy (tfht):

where did you get h from? is it the center?

OpenStudy (anonymous):

yes

OpenStudy (tfht):

ok then you would do dh=fh and gh=eh. so it would look like- x + 5=2y and 3x – 1=5y + 4 then you would set it up so that x= whatever in a problem. so it could be x=2y-5 and sub that in for x in the other equation to get 3(2y-5)-1=5y+4

OpenStudy (tfht):

then get 6y-16=5y+4 then get y alone and you get y=20

OpenStudy (tfht):

then sub in 20 for y to get x

OpenStudy (tfht):

so you would get x+5=(20)2 so you would get x=40-5 and so x=35